My hundreds digit is 4 less than my tens digit. my ten thousands digit is 6 greater than my hundreds digit. my thousand digit is 5 less than my thousand digit and my ones digit is 6 greater than my hundred digit. if my tens digit is 7, what number am i ?
Found 2 solutions by Aldorozos, ikleyn:
Answer by Aldorozos(172)


There is an issue here where we say my (thousand digit is 5 less than my thousand digit). How about if we reword that and say (my thousand digit is 5 less than my ten thousand digit instead of my 1000 digit.) My guess is that thousand digit can not be 5 less than my thousand digit)
For that reason I rewrite the puzzle
My hundreds digit is 4 less than my tens digit. my ten thousands digit is 6 greater than my hundreds digit. my thousand digit is 5 less than my hundred digit and my ones digit is 6 greater than my hundred digit. if my tens digit is 7, what number am i ?
Lets assume the number is
nmxyz where n is 10000 digit, m is 1000 digit x is 1000 digit, y is 10 digit and z is ones digit
Then we can wtire
x=y-4
n=x+6
m=n-5
z=x+6
if y =4 what are x, n, m and z
if y = 4 then x=y-4 = 4-4 = 0
if x is zero the n = x+6 = 0+6
if n = 6 then m = 6-5 = 1
if x = 0 then z = 0+6

Answer by ikleyn(52756)

.
The condition is wrong regarding the part related to the "thousand digit".